# Multi-Agent Retrieval-Augmented Framework for Evidence Based Counterspeech Against Health Misinformation

### v1

@@ 4.2 Baseline
LLM With Direct Prompt (LLM w DP) This method uses direct instructions (Appendix  A.4.1) to prompt LLM to generate counterspeech without any external knowledge. The model takes health misinformation posts as input and follows direct instructions to generate responses using LLM’s internal knowledge.
LLM With Prompt Engineering (LLM w PE) This method uses the refined Guided Instruction prompt (Appendix A.4.2) to structure the LLM’s responses, ensuring clarity, factual accuracy, and coherence, however, still relying on the LLM’s internal knowledge.
Static RAG This approach employs RAG to extract supporting evidence from our curated knowledge bases, enabling us to validate the effectiveness of the static evidence in generating accurate and evidence-based responses. The retrieved evidence is then integrated into a Guided Instruction prompt (Appendix A.4.2), guiding the LLM in generating evidence-based counterspeech.
Dynamic RAG This approach retrieves real-time evidence from the DuckDuckGo search engine, using the misinformation post as a query. By leveraging real-time web search, this method enables dynamic retrieval of the most recent and relevant information, allowing for a comparative analysis between dynamic evidence and static knowledge bases. Notably, this retrieval process follows the same Guided Instruction prompt framework, allowing a comparative assessment of the effectiveness of dynamic evidence in generating counterspeech.
@@ 4.3 Multi-Agent Retrieval-Augmented Framework Implementation
The Multi-agent Retrieval-Augmented Framework consists of multiple specialized agents that operate sequentially to optimize the retrieval, filtering, summarization, generation, and refinement of responses to health misinformation. We use the ”Llama-3.1-8B-Instruct” (Grattafiori et al., 2024) model for generation. Generation configurations are detailed in Appendix A.3.
@@
Knowledge Source In terms of the static knowledge source, we collect evidence from trusted sources such as the National Institutes of Health (NIH)  [^6] https://www.ncbi.nlm.nih.gov  [^7] https://covid19community.nih.gov World Health Organization (WHO)  [^8] https://www.who.int  [^9] https://www.who.int/docs/default-source  [^10] https://www.who.int/emergencies/diseases. These institutions are globally acknowledged for their authoritative presence and credibility, offering regular updates on health-related information that are crucial for creating trustworthy counterspeech (Kington et al., 2021). Our dynamic knowledge source utilizes the DuckDuckGo search engine API to retrieve real-time, web-based information. This method enables the system to access current evidence from credible news outlets, government health websites, and fact-checking organizations.
@@
Static Retrieval Agent This Agent performed a Hybrid search in our curated knowledge base using both Vector-Based Retrieval and Keyword-Based Retrieval. Vector-based retrieval employs a pre-trained embedding model ”sentence-transformers/all-mpnet-base-v2”  [^11] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for embedding-based similarity searches and ranking by cosine similarity, while Keyword-Based Retrieval utilizes ”SimpleKeywordTableIndex” to identify exact team matches within the documents.
@@
Dynamic Retrieval Agent The Dynamic Retrieval Agent uses the DuckDuckGo API [^12] https://pypi.org/project/duckduckgo-search-api/ to fetch live external evidence. The misinformation post serves as the structured search query. It extracts key factual sentences from each result via web scraping (using BeautifulSoup [^13] https://pypi.org/project/beautifulsoup4/). By querying DuckDuckGo, the Dynamic Retrieval Agent can capture the latest posts, news articles, and trusted websites relevant to a given health misinformation post. Additionally, we conduct fact-checking and filter the information with lower factual accuracy, ensuring that the selected information maintains an average accuracy score above 0.65.
@@
Summarization Agent This Agent ranks the retrieved evidence collected from the Static Retrieval Agent and the Dynamic Retrieval Agent. We evaluate various top-k evidence selections, as illustrated in Figure 2, which indicates that top-10 evidence yields a slightly better performance across all dimensions. Therefore, we filter the top-10 evidence using BM25-based ranking and semantic embedding similarity and summarize them using the prompt in Appendix  A.4.3, effectively preserving key factual details.
@@
Counterspeech Generation Agent The summarized evidence is then further utilized by the LLM to generate fact-based counterspeech. This agent employs a structured prompting approach (Appendix  A.4.2), which effectively guides the LLM to produce concise, informative, and contextually relevant responses.
@@
Refinement Agent The final step involves refining the generated responses using the Refinement Agent. A refined prompt (Appendix A.4.4) guides this process, ensuring that responses are respectful, engaging, and persuasive.
@@
Table 1: Comparison of Multi-Agent counterspeech generation with baseline methods across key evaluation metrics. (DP: direct prompt, and PE: prompt engineering.) Higher values indicate better performance. The best scores in each category are underlined.
@@ 4.4 Evaluation Results
Table 1 presents the results of the experiment. The overall performance analysis indicates that LLM w DP exhibits slightly strong relevance (0.70) and informativeness (0.77), compared with other baseline methods. Prompt engineering (LLM w PE) significantly improves politeness (0.84) and factual accuracy (0.83) compared with LLM w DP. However, it comes at the cost of slightly lower relevance (0.65).
Introducing static knowledge retrieval (Static RAG) provides better grounding by integrating factual information from a structured knowledge base. However, the results indicate that this does not result in a significant overall improvement across all dimensions compared with LLM without external knowledge (LLM w DP and LLM w PE), suggesting that LLM may struggle to effectively incorporate external knowledge into generation (Bondarenko & Viehweger, 2024). This pattern remains consistent in the Dynamic RAG experiment (see Figure 3), even when real-time external evidence is provided, further validating that LLMs may have inherent limitations in effectively integrating external knowledge.
Our Multi-agent Retrieval-Augmented Framework successfully addresses these gaps by effectively integrating multiple retrieved evidence and refining the final responses. As a result, it achieves higher politeness (0.88), informativeness (0.78), and factual accuracy (0.86) while maintaining substantial relevance (0.70), making it a more effective and balanced approach to countering misinformation.
@@ 4.5 Ablation Experiments
To demonstrate the necessity of multiple agents in our framework, we conduct ablation experiments to generate counterspeech. Additionally, we explore various prompting strategies to assess whether they lead to performance improvements. Table 2 represents how the involvement of different agents contributes to the performance of counterspeech generation under the Guided and CoT prompting strategies (See in Appendix A.4).
MA w/o SA w/o RF shows the effect after the removal of Summarization and Refinement Agents. This setup only relies on retrieval and response generation without any additional refinement. It maintains relatively high scores for factual accuracy (Guided: 0.85) and informativeness (Guided: 0.78), however, with lower politeness (Guided: 0.80) compared with MA. Including the Summarization Agent (MA w/o RF with Guided) improves the relevance (0.70), factual accuracy (0.87), and informativeness (0.82) compared with MA w/o SA w/o RF. This further demonstrates that summarized evidence helps LLMs better integrate evidence into generation. Furthermore, the addition of the Refinement Agent, our proposed Multi-Agent approach (MA (Ours)), further refines responses by ensuring they are organized, polite, and clear, with the best performance across all dimensions.
We also examine the effect of changing the Prompting Strategy to Chain-of-Thought (CoT) (See details in Appendix A.4.5). While the CoT prompting strategy enables the generation to maintain similar factual accuracy compared with Guided prompting, the performance of politeness (MA w/o SA w/o RF: 0.63 vs 0.80; MA w/o RF: 0.65 vs 0.79; MA: 0.74 vs 0.88), and informativeness (MA w/o SA w/o RF: 0.76 vs 0.78; MA w/o RF: 0.77 vs 0.82; MA: 0.77 vs 0.78) decrease. The results indicate that CoT reasoning may be less effective than Guided prompting in our study. The step-by-step reasoning process often introduces redundant or overly detailed content, which decreases politeness and informativeness (Sprague et al., 2024; Wei et al., 2022).
@@
Table 2: Evaluation of the impact of Summarization Agent (SA) and Refinement Agent (RF) alongside different prompting styles (Guided vs. CoT) on Multi-Agent (MA) counterspeech generation performance.
@@ 5 Human Evaluation
We conduct two rounds of human evaluations. The first round is to validate the LLM evaluation. Even though various studies rely on LLMs as evaluators(Li et al., 2024b; Evans et al., 2024; Chen et al., 2024), different research tasks can yield varying LLM performance (Chern et al., 2024; Chan et al., 2023). Therefore, we involve humans to verify LLM evaluation of informativeness and factual accuracy. The second round aims to explore the effectiveness of the refined response. In our study, we utilize the Refinement agent to improve counterspeech, however, it remains unknown whether users prefer these refined responses. Therefore, it is necessary to figure out human preference, which may help us improve generation in future studies.
In the first stage, we conduct a human evaluation to verify the reliability of the computation evaluation: Informativeness and Factual Accuracy. The primary objective is to determine how well the model-generated evaluation scores align with human judgments. We randomly select 100 samples, and three annotators independently evaluate the generated responses based on informativeness and factual accuracy on a 1-5 scale (Details guidelines in Appendix A.5). Initially, the broad rating scale results in variability in annotator agreement. To address this, further discussions among the annotators are held, leading to a final label that reflects a consensus-driven evaluation. Once the final human labels are established, we compared them with our model evaluation results. For factual accuracy, Cohen’s Kappa is $\kappa\geq 0.69$, and for informativeness, Cohen’s Kappa is $\kappa\geq 0.65$, indicating substantial agreement with the model evaluation. These results suggest that the computing-based evaluation exhibits strong consistency with human judgments.
Secondly, we assess the effectiveness of the Refinement Agent in counterspeech responses. We conduct a comparative human evaluation of two different response strategies: Response A (Multi-Agent w/o Refinement Agent) and Response B (Multi-Agent w Refinement Agent). The goal is to determine human preference and assess how the Refinement Agent enhances response quality(details of the guidelines in Appendix A.5). The initial agreement rates among them are 74%, 71%, and 73%, respectively, which shows they have a relatively high level of consistency. The disagreements were solved by further discussion and concluded with a final label. The annotation results reveal that 72% of users prefer Response B, which deploys a Refinement Agent, indicating that clearer and better-organized responses are more accessible for users. In the future, we can continue refining our methods to produce counterspeech that is even more user-friendly and effective.
@@ 6 CONCLUSION
In this study, we present a Multi-Agent Retrieval-Augmented Framework for generating effective counterspeech against health misinformation. This framework integrates multiple specialized LLM agents to sequentially (1) retrieve evidence from both static knowledge bases and dynamic web sources, (2) filter and summarize the combined evidence, and (3) generate and refine counterspeech using structured prompting. This design enables the generation of responses that are not only factually accurate but also polite, informative, and contextually relevant. Experimental results and human evaluations demonstrate that our method outperforms several baselines across multiple dimensions. Additionally, ablation studies confirm the importance of each module, particularly the refinement agent, which plays a key role in enhancing the clarity and engagement of counterspeech. Our findings highlight the value of combining static and dynamic knowledge and the refined method to improve the reliability and quality of automated counterspeech. We acknowledge several limitations in our study. Even though we incorporate both static and dynamic evidence, we recognize that we have not fully collected all existing static knowledge since the cost of manual work. Moreover, there may be some overlap between the two knowledge sources. Nevertheless, we claim that each knowledge base serves a distinct purpose: static evidence offers rigorously validated information, while dynamic evidence provides real-time updates. And the proposed Multi-Agent Retrieval-Augmented Framework enables efficient integration of these components to generate effective counterspeech. In future work, we aim to further refine the curation processes for both knowledge sources and explore methods to minimize overlap, enhancing the accuracy of our framework in addressing health misinformation.
@@
Our study follows the highest ethical standards to address misinformation using Large language models. We acknowledge that AI-generated responses can affect society, and we are careful to ensure they are effective while still being ethical. We use publicly available data sources, ensuring that no personally identifiable information is collected or used. User identities from social media datasets are anonymized, and any sensitive information is removed before analysis. Human annotators involved in data annotation and evaluating counterspeech responses remain anonymous, and their participation is voluntary. We follow rules to make sure the annotation process is fair and open. Our study aims to generate respectful, fact-based counterspeech while minimizing unintended biases. We understand that LLMs can show biases based on the data on which they were trained. To address this, we use retrieval-augmented methods to ensure our answers are based on verified facts. We ensure that the AI-generated responses are not misleading, confrontational, or harmful. The system is created to give polite and well-structured responses that encourage helpful discussion rather than increasing division. We are committed to transparency in our methodology, which includes providing details on model architectures, data sources, and evaluation criteria. Our work aligns with the principles of ethical AI research and aims to positively contribute to public discourse by countering misinformation with reliable evidence. By following these ethical considerations, We aim to create AI solutions that encourage good communication, build public trust, and deliver accurate, respectful, and reliable information.

### v2

@@ 4.2 Baseline
LLM with Direct Prompt (LLM w DP) This method used direct instructions (Appendix  A.4.1) to prompt the LLM to generate counterspeech without any external knowledge. The model took health misinformation posts as input and followed direct instructions to generate responses using LLM’s internal knowledge.
LLM with Prompt Engineering (LLM w PE) This method used the refined Guided Instruction prompt (Appendix A.4.2) to structure the LLM’s responses, ensured clarity, factual accuracy, and coherence while still relying on the LLM’s internal knowledge.
Static RAG This approach employed RAG to extract supporting evidence from our curated knowledge bases, enabling us to validate the effectiveness of the static evidence in generating accurate and evidence-based responses. The retrieved evidence was then integrated into a Guided Instruction prompt (Appendix A.4.2), guiding the LLM in generating evidence-based counterspeech.
Dynamic RAG This approach retrieved real-time evidence from the DuckDuckGo search engine, using the misinformation post as a query. By leveraging real-time web search, this method enables dynamic retrieval of the most recent and relevant information, allowing for a comparative analysis between dynamic evidence and static knowledge bases. Notably, this retrieval process followed the same Guided Instruction prompt, allowing a comparative assessment of the effectiveness of dynamic evidence in generating counterspeech.
@@ 4.3 Multi-Agent Retrieval-Augmented Framework Implementation
The Multi-agent Retrieval-Augmented Framework consists of multiple specialized agents that operate sequentially to optimize the retrieval, filtering, summarization, generation, and refinement of responses to health misinformation. We used the ”Llama-3.1-8B-Instruct” (Grattafiori et al., 2024) model for generation. Generation configurations are detailed in Appendix A.3.
@@
Knowledge Source In terms of the static knowledge source, we collected evidence from trusted sources such as the National Institutes of Health (NIH)  [^7] https://www.ncbi.nlm.nih.gov/  [^8] https://covid19community.nih.gov/, World Health Organization (WHO)  [^9] WHO Misinformation Toolkit  [^10] WHO Roles & Responsibilities  [^11] WHO FAQ on COVID-19. These institutions are globally acknowledged for their authoritative presence and credibility, offering regular updates on health-related information that are crucial for creating trustworthy counterspeech (Kington et al., 2021). Our dynamic knowledge source utilized the DuckDuckGo search engine API to retrieve real-time, web-based information. This method enables the system to access current evidence from credible news outlets, government health websites, and fact-checking organizations. To empirically assess the potential overlap between the two sources, we computed the BLEU score (Papineni et al., 2002) between retrieved static and dynamic evidence across our samples. The results show a mean BLEU score of 0.0052 with a standard deviation of 0.0033, indicating minimal lexical overlap. This suggests that while dynamic knowledge may occasionally cite static sources, it generally provides distinct and complementary information, supporting the separation used in our system design.
@@
Static Retrieval Agent This Agent performed a hybrid search in our curated knowledge base using both Vector-Based Retrieval and Keyword-Based Retrieval. Vector-based retrieval employed a pre-trained embedding model ”sentence-transformers/all-mpnet-base-v2”  [^12] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for embedding-based similarity searches and ranking by cosine similarity, while Keyword-Based Retrieval utilized ”SimpleKeywordTableIndex” to identify exact team matches within the documents.
@@
Dynamic Retrieval Agent The Dynamic Retrieval Agent useed the DuckDuckGo API [^13] https://pypi.org/project/duckduckgo-search-api/ to fetch live external evidence. The misinformation post serves as the structured search query. It extracted key factual sentences from each result via web scraping (using BeautifulSoup [^14] https://pypi.org/project/beautifulsoup4/). By querying DuckDuckGo, the Dynamic Retrieval Agent can capture the latest posts, news articles, and trusted websites relevant to a given health misinformation post. Additionally, we conducted fact-checking and filtered the information with lower factual accuracy, ensuring that the selected information maintains an average accuracy score above 0.65.
@@
Summarization Agent This Agent ranked the retrieved evidence collected from the Static Retrieval Agent and the Dynamic Retrieval Agent. We evaluated various top-k evidence selections, as illustrated in Figure 2, which indicates that top-10 evidence yields a slightly better performance across all dimensions. Therefore, we filtered the top 10 evidence using BM25-based ranking and semantic embedding similarity, and summarized them using the prompt in Appendix  A.4.3, effectively preserving key factual details.
@@
Counterspeech Generation Agent The summarized evidence was then further utilized by the LLM to generate fact-based counterspeech. This agent employed a structured prompting approach (Appendix  A.4.2), which effectively guides the LLM to produce concise, informative, and contextually relevant responses.
@@
Refinement Agent The final step involved refining the generated responses using the Refinement Agent. A refined prompt (Appendix A.4.4) guided this process, ensuring that responses are respectful, engaging, and persuasive.
@@
Table 1: Comparison of Multi-Agent counterspeech generation with baseline methods across key evaluation metrics. (DP: direct prompt, and PE: prompt engineering). Higher values indicate better performance. The best scores in each category are underlined. Each score is reported as mean (standard deviation).
@@ 4.4 Evaluation Results
Table 1 presents the results of the experiment. The overall performance analysis indicated that LLM w DP exhibited slightly strong relevance (0.70) and informativeness (0.77), compared with other baseline methods. Prompt engineering (LLM w PE) significantly improved politeness (0.84) and factual accuracy (0.83) compared with LLM w DP. However, it comes at the cost of slightly lower relevance (0.65).
Introducing static knowledge retrieval (Static RAG) provided better grounding by integrating factual information from a structured knowledge base. However, the results indicated that this did not result in a significant overall improvement across all dimensions compared with LLM without external knowledge (LLM w DP and LLM w PE), suggesting that LLM may struggle to effectively incorporate external knowledge into generation (Bondarenko & Viehweger, 2024). This pattern remained consistent in the Dynamic RAG experiment (see Figure 3), even when real-time external evidence was provided, further validating that LLMs may have inherent limitations in effectively integrating external knowledge.
Our Multi-agent Retrieval-Augmented Framework successfully addressed these gaps by effectively integrating multiple retrieved evidence and refining the final responses. As a result, it achieved higher politeness (0.88), informativeness (0.78), and factual accuracy (0.86) while maintaining substantial relevance (0.70), making it a more effective and balanced approach to countering misinformation.
@@ 4.5 Ablation Experiments
To demonstrate the necessity of multiple agents in our framework, we conducted ablation experiments to generate counterspeech. Additionally, we explored various prompting strategies to assess whether they lead to performance improvements. Table 2 represents how the involvement of different agents contributed to the performance of counterspeech generation under the Guided and CoT prompting strategies (See in Appendix A.4).
@@
Table 2: Evaluation of the impact of Summarization Agent (SA) and Refinement Agent (RF) alongside different prompting styles (Guided vs. CoT) on Multi-Agent (MA) counterspeech generation performance. Each score is reported as mean (standard deviation).
MA w/o SA w/o RF showed the effect after the removal of Summarization and Refinement Agents. This setup only relied on retrieval and response generation without any additional refinement. It maintained relatively high scores for factual accuracy (Guided: 0.85) and informativeness (Guided: 0.78), however, with lower politeness (Guided: 0.80) compared with MA. Including the Summarization Agent (MA w/o RF with Guided) improved the relevance (0.70), factual accuracy (0.87), and informativeness (0.82) compared with MA w/o SA w/o RF. This further demonstrated that summarized evidence helped LLMs better integrate evidence into generation. Furthermore, the addition of the Refinement Agent, our proposed Multi-Agent approach (MA (Ours)), further refined responses by ensuring they were organized, polite, and clear, with the best performance across all dimensions.
We also examined the effect of changing the Prompting Strategy to Chain-of-Thought (CoT) (See details in Appendix A.4.5). While the CoT prompting strategy enabled the generation to maintained similar factual accuracy compared with Guided prompting, the performance of politeness (MA w/o SA w/o RF: 0.63 vs 0.80; MA w/o RF: 0.65 vs 0.79; MA: 0.74 vs 0.88), and informativeness (MA w/o SA w/o RF: 0.76 vs 0.78; MA w/o RF: 0.77 vs 0.82; MA: 0.77 vs 0.78) decreased. The results indicated that CoT reasoning may be less effective than Guided prompting in our study. The step-by-step reasoning process often introduced redundant or overly detailed content, which decreased politeness and informativeness (Sprague et al., 2025; Wei et al., 2022).
4.6 Cross Evaluation
Topic-wise Evaluation To further assess the robustness of our framework across different types of misinformation, we conducted a topic-wise analysis covering COVID-19, HIV, and Influenza. As shown in Appendix A.5 (Table 7), our Multi-Agent framework consistently achieved strong average scores across all topics and evaluation dimensions, with the following scores: politeness (0.90), relevance (0.71), informativeness (0.77), and factual accuracy (0.85). This further demonstrated its generalizability across diverse health misinformation domains.
Method Politeness Relevance Informativeness Factual Accuracy
LLM w DP 0.44 (0.26) 0.70 (0.14) 0.77 (0.11) 0.81 (0.21)
LLM w PE 0.81 (0.18) 0.77 (0.11) 0.77 (0.07) 0.94 (0.12)
Static RAG 0.86 (0.14) 0.77 (0.12) 0.78 (0.09) 0.95 (0.12)
Dynamic RAG 0.85 (0.16) 0.77 (0.13) 0.76 (0.10) 0.92 (0.16)
Multi-Agent (Ours) 0.89 (0.11) 0.78 (0.10) 0.83 (0.12) 0.96 (0.12)
Table 3: Cross-platform generalization results on the MisinfoCorrect dataset (He et al., 2023). Our method achieves the highest performance across all evaluation metrics. (DP: direct prompt, and PE: prompt engineering). Each score is reported as mean (standard deviation).
Cross-Platform Generalization We evaluated our system on the external MisinfoCorrect dataset (He et al., 2023), which contains COVID-19 vaccine misinformation from Twitter (now X). This experiment tests cross-platform generalization. Results, shown in Table 3, demonstrated that our method maintained top performance on this cross-platform dataset, with politeness (0.89), relevance (0.78), informativeness (0.83), and factual accuracy (0.96).
@@ 5 Human Evaluation
We conducted multiple rounds of human evaluations to assess the quality and effectiveness of the generated counterspeech comprehensively. The first round was to validate the LLM evaluation. Even though various studies relied on LLMs as evaluators (Li et al., 2024b; Evans et al., 2024; Chen et al., 2024), different research tasks could yield varying LLM performance (Chern et al., 2024; Chan et al., 2024). Therefore, we involved humans to verify the LLM’s evaluation of informativeness and factual accuracy. The second round aimed to explore the effectiveness of the refined response. In our study, we utilized the Refinement Agent to improve counterspeech; however, it remained unknown whether users preferred these refined responses. Therefore, it was necessary to determine human preference, which might help us improve generation in future studies. In the final round of human evaluation, we assessed the overall quality of counterspeech across systems. We compared responses generated by our proposed Multi-Agent framework against two strong baselines, Static RAG and Dynamic RAG, based on human judgments of effectiveness in correcting health misinformation.
In the first stage, we conducted a human evaluation to verify the reliability of the computational evaluation, focusing on informativeness and factual accuracy (details are provided in Appendix A.6). For factual accuracy, Cohen’s Kappa was $\kappa\geq 0.69$, and for informativeness, Cohen’s Kappa was $\kappa\geq 0.65$, which indicated substantial agreement with the model evaluation. These results suggested that the computing-based evaluation exhibited strong consistency with human judgments.
Secondly, we assessed the effectiveness of the Refinement Agent in counterspeech responses (details of the guidelines in Appendix A.7). The initial agreement rates among them were 74%, 71%, and 73%, respectively, which showed they had a relatively high level of consistency. The annotation results showed that 72% of users preferred the responses generated with the Refinement Agent, which indicated that clearer and better-organized responses are more accessible for users. In the future, we could continue refining our methods to produce counterspeech that is even more user-friendly and effective.
In the final round, we compared the overall effectiveness of our proposed Multi-Agent Framework with two strong baselines, Static RAG and Dynamic RAG (detailed in Appendix A.8). The inter-annotator agreement was substantial, with an agreement rate of 80% and a Cohen’s Kappa of $\kappa\geq 0.69$. Results showed that the counterspeech generated by our proposed framework was selected as the most effective in 47% of cases, compared to 32% for Static RAG and 21% for Dynamic RAG. These findings suggested that our method achieved superior overall quality based on human evaluations.
@@ 6 CONCLUSION
In this study, we presented a Multi-Agent Retrieval-Augmented Framework for generating effective counterspeech against health misinformation. This framework integrated multiple specialized LLM agents to sequentially (1) retrieve evidence from both static knowledge bases and dynamic web sources, (2) filter and summarize the combined evidence, and (3) generate and refine counterspeech using structured prompting. This design enables the generation of responses that were not only factually accurate but also polite, informative, and contextually relevant. Experimental results and human evaluations demonstrated that our method outperforms several baselines across multiple dimensions. Additionally, ablation studies confirmed the importance of each module, particularly the refinement agent, which played a key role in enhancing the clarity and engagement of counterspeech. Our findings highlighted the value of combining static and dynamic knowledge and the refined method to improve the reliability and quality of automated counterspeech. We acknowledge several limitations in our study. We recognize that we haven’t fully collected all existing static knowledge due to the cost of manual work, and our system may require many computing resources. In future work, we aim to further refine the curation processes for knowledge sources and explore methods to improve the efficiency of our framework.
@@
Our study follows the highest ethical standards to address misinformation using Large Language Models. We acknowledge that AI-generated responses can affect society, and we are careful to ensure they are effective while still being ethical. We use publicly available data sources, ensuring that no personally identifiable information is collected or used. User identities from social media datasets are anonymized, and any sensitive information is removed before analysis. Human annotators involved in data annotation and evaluating counterspeech responses remain anonymous, and their participation is voluntary. Annotators are compensated on average with \$15 per hour. We follow rules to make sure the annotation process is fair and open. Our study aims to generate respectful, fact-based counterspeech while minimizing unintended biases. We understand that LLMs can show biases based on the data on which they were trained. To address this, we use retrieval-augmented methods to ensure our answers are based on verified facts. We ensure that the AI-generated responses are not misleading, confrontational, or harmful. The system is created to give polite and well-structured responses that encourage helpful discussion rather than increasing division. We are committed to transparency in our methodology, which includes providing details on model architectures, data sources, and evaluation criteria. Our work aligns with the principles of ethical AI research and aims to positively contribute to public discourse by countering misinformation with reliable evidence. By following these ethical considerations, we aim to create AI solutions that encourage good communication, build public trust, and deliver accurate, respectful, and reliable information.